If (1+ tan 1) (1+tan 2)…..(1+tan45) = 2^n , then the value of n is…??
(a). 21
(b). 22
(c). 23
(d). 24
Grade:

4 Answers

Chetan Mandayam Nayakar
12 years ago

Perhaps, the formula tan((π/4)+θ)=(1+tan(theta))/(1-tan(theta)) may help.

Hope it works

Swapnil Saxena
12 years ago

The answer is 23

(1+ tan1) = 1+tan(45-44) = 1+ (tan(45) - tan(44))/1+tan(45)tan(44)  = 1+ ((1-tan44)/(1+tan44)) ==> 2/(1+tan(44))

Thus the above qproblem can be rewritten as

(2/(1+tan(44)))(2/(1+tan(43)))...(2/(1+tan(23))) * (1+tan(23))(1+tan(24))(1+tan(25))...(1+tan(45))

Can be simplified as (2*2*2...22 times)*(2) = 2^23

Thus the value of n is 23

vinay chaudhary
12 years ago

answer is 23.

since,(1+tan1)can be written as (1+tan(45-44)

on solving ,

we get (1+(1-tan44)/(1+tan44))

then , multiply the second last term 

(1+(1-tan44)/(1+tan44))*(1+tan44)=2

similarly, all other terms can be solved.

Vaishal Shah
12 years ago
if A+B=45 then (1+tanA)(1+tanB)=2.. So,the ans. Is 23.